the value of p and q such that f(x)=x4-px2+q leaves remainder 3x+4 when divided by x2-3x+2
 
 
 

Grade:11

1 Answers

Vikas TU
4 years ago
Hiii 
X^2-3x + 2 can be written as X^2-2x-x+2
=> (x-2)(x-1)
Divident = Divisor * Quotient + Remainder 
X^4 – px^2 +q = (X^2-3x + 2 ) +  3x+4
Put x=1
p-q = 6 …...(1)
Put x=2 
-4p+q = -6......(2)
1 + 2 we get 
-3p = -12 
p= 4
Put in 1 
q=10